# Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge

### v8

@@ Abstract
Lin Shi, Chiyu Ma, Wenhua Liang, Xingjian Diao, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, xingjian.diao.gr, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability in extensive evaluations Zeng et al. (2023); Karpinska et al. (2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments Zheng et al. (2024b); Li et al. (2024a); Zhu et al. (2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1 (a). This bias has been prevalent in all types of LLM judges Qin et al. (2024); Li et al. (2023d), raising concerns about their reliability. Previous studies Zheng et al. (2024a, b); Zeng et al. (2023) have identified position bias alongside other biases and assessed its impact. Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al. (2024), increased complexity Li et al. (2024b); Khan et al. (2024); Chua et al. (2024), the introduction of new biases Ohi et al. (2024), inconsistent effectiveness Gallegos et al. (2024), or impracticality for closed-source models, as mechanistic solutions require modifications to model internals Wang et al. (2025); Yu et al. (2025). From an exploratory perspectives, despite empirical studies evaluating position bias in pairwise LLM-as-a-Judge settings Zheng et al. (2024b); Wang et al. (2023), the underlying factors contributing to this phenomenon and extended measurements in list-wise comparative paradigms remain incompletely explored.
In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge. While evaluating Position Consistency Zheng et al. (2024b), one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability. Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences. Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
@@ 2.2 Evaluation Metrics
Preference Fairness ($PF$) measures the extent to which LLM judges favor certain solution positions. In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference/bias for the first/second candidate model" used in previous studies Zheng et al. (2024b), ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 1 (b). Previous studies proposed two common ways to measure the preference fairness. One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach overlooks the fact that “position consistent judgments are also preference fair”, which leads to overly penalizing highly consistent LLM-judges.
@@
$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
@@ 2.3 Factors Affecting Position Bias
Answer quality gap: While prior studies Wang et al. (2023) explored quality disparities using "score gaps" in score-based LLM-as-a-Judge, this factor remains under-explored in comparative settings, which we address by introducing "answer quality gap" for both pairwise and list-wise evaluations. We define the quality of a candidate’s solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the LLM judge.
Following this assumption, we measure the answer quality gap by the win rates of candidates over an expected baseline on a set of tasks and questions. However, if position bias occurs, the winner may be inconsistent when the order of candidate solutions is permuted in the query. Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen across all permutations (termed “consistent wins”), cases where there is no certain winner (termed “consistent ties”), and cases where different winners are selected after the solutions are permuted in the queries (termed “inconsistent judgment series"). We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment series ($C_{I}$), respectively. Inspired by Zheng et al., we count inconsistent judgment pairs as ties for all candidate models, which is later calculated as a down-scaled win rate depending on the number of candidate models.
To calculate the win rates of candidate models for all three cases, we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{p}(C_{t}+C_{I})}]$, where we have $p$ candidates in the list and $n$ judgment series. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-1/p|$, where $1/p$ is the expected baseline when all judgments are “ties”. In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment series) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
@@ 3.1 Experiment Settings
In this study, we evaluated position bias of 15 models from the GPT (OpenAI, 2023), Claude (Anthropic, 2024), Gemini (Gemini Team, 2024), and Llama (Touvron et al., 2023) series using our framework. For datasets, we adopted the modified MTBench (Zheng et al., 2024b) and DevBench (Li et al., 2024a) due to their demonstrated high human-LLM agreement and the validated reliability of state-of-the-art LLM judges on the evaluation tasks. For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of solutions to the given questions. MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges.
@@ 3.2 Empirical Results
The evaluation results of 12 close-source and 3 open-source LLM judges in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, $PC$, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not due to random variations. Fig. 2 (a)(b) demonstrate that position bias varies by judges and tasks significantly. Fig. 2 (c) explores the correlation between the metrics $PC$ and $PF$. Fig. (d) to (f) further investigate the impact of the answer quality gap on position bias. These analyses were conducted by considering all judges together on MTBench. More analyses can be found in Appendix. Sec. D.
@@
$RS$ $PC$ $PF$ Error $RS$ $PC$ $PF$ Error $PC$ $PF$ Error
@@
Table 2: Evaluation results for Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold. Errors arise from judgment failures (e.g., exceeding context window, not following output format). High error rates and low $RS$ are marked red, rendering further evaluations invalid due to insufficient data. List-wise evaluation is conducted on four representative judges to validate scalability.
@@ 4 Main Findings
Figure 2: Judge performances on MTBench. Fig. (a)(b) are the radar charts for the $PC$ comparison by family, judge, and task. Fig. (c) leverages linear regression to explore the general relationship between $PC$ and $PF$. Fig. (d) to (f) investigate the impact of answer quality gap on position bias using overall win rates. Fig. (a) to (e) are for pairwise comparative settings, while Fig. (f) are obtained under list-wise evaluations.
@@
Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate ($owr$). Therefore, $\delta_{q}$ increases as $owr$ extends from baseline to 0 or 1. Fig. 2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs or series with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. D.
@@ B.1 LLM-as-a-Judge
Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments. The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (Chiang and Lee, 2023a), adversarial attacks (Chiang and Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi and Federmann, 2023), and instruction following (Zeng et al., 2023), where models are tasked with scoring or ranking outputs. Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness, even in the most effective pairwise comparative settings.
@@ B.2 Position Bias
However, position bias is a complex problem that is challenging to deal with or solve. The naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited. Therefore, researchers proposed ‘inconsistency-as-a-tie’ for both models in pairwise comparative settings to consider all judgments for further analysis (Zheng et al., 2024b; Li et al., 2023c). This approach, while practically useful for evaluations, does not mitigate position bias.
Given the significance of position bias, researchers have developed more sophisticated approaches to mitigate it, including bootstrapping (Hou et al., 2024), split-and-merge techniques (Li et al., 2023d), and multi-agent discussions (Li et al., 2023c; Khan et al., 2024). While these methods demonstrate potential, they are often costly, time-consuming, or insufficient. For example, multi-agent discussions and iterative reviews require extensive computational resources and human-like deliberation, while simpler techniques, such as order permutation (Wang et al., 2023; Zheng et al., 2024a) and batch calibration (Zhou et al., 2024), have limited effectiveness in addressing bias holistically. Furthermore, calibration methods like ensembling (Li et al., 2023a) or context decomposition (Li et al., 2023d) are often constrained by the inadequacy of supervised data to align LLM evaluators effectively (Liu et al., 2024). This highlights that addressing this crucial bias is complex and challenging.
The limitations of existing approaches underscore a critical gap in our understanding of position bias itself. A lack of clarity on the factors influencing this bias and their quantitative impacts hampers the development of effective solutions. For instance, while the PORTIA approach (Li et al., 2023d) demonstrates significant improvements in consistency for GPT-4 (from 93.44% to 97.03%), this success may be overestimated due to the selection of tasks with large answer quality gaps. As our study reveals, such gaps inherently reduce the likelihood of position bias, making these cases easier to calibrate. This example underscores the necessity of first comprehensively understanding the underlying factors contributing to position bias before proposing or refining mitigation strategies. Without this foundational understanding, the efficacy of current and future methods remains uncertain and potentially misjudged.
B.3 Research Gaps
To enhance the understanding of position bias, our study builds on and extends prior work through the following dimensions.
Repetition Stability: Previous studies on LLM-as-a-Judge often implicitly assume that single-instance judgments are sufficient to reflect the evaluating capabilities of judge models, thereby attributing observed biases entirely to systematic tendencies rather than random variations. These studies typically proceed directly to analyses on position bias or others without such verification. However, if judgments vary by chance or lack consistency across repetitions, the reliability and accuracy of downstream analyses are compromised. To avoid this issue, Chen et al. (2024) repeated evaluations multiple times and adopted the mode judgment as the representative, an approach that, while effective, is computationally expensive. Addressing this gap, our work introduces a method or metric for systematically evaluating the stability of LLM judgments across repetitions. This ensures that observed biases are systematic rather than artifacts of variability. By performing this validation on a sample dataset for cost-efficiency, researchers can establish a robust foundation for subsequent evaluations and analyses, ensuring the reliability of their conclusions.
Preference Fairness: Prior studies on position bias often focus on the proportion of position consistent judgments, referred to as “position consistency”. Only a few have explored the preference directions of LLM evaluators in addition to consistency. However, traditional measurements either overly penalize highly consistent judges (Li et al., 2023c; Liusie et al., 2024) or rely on metrics whose sensitivity is highly dependent on dataset size (Zheng et al., 2024b; Zhu et al., 2023). To address these limitations, we proposed a refined metric, preference fairness ($PF$), which integrates the strengths of both approaches. This single, comprehensive metric ensures sensitivity across datasets, remains independent of the number of questions and tasks, and represents a significant improvement over previous metrics.
Answer Quality Gap: Intuitively, candidate solutions with significant quality disparities are easier to judge, likely resulting in reduced position bias. This suggests that the quality gap among answers is a potentially significant factor influencing position bias. However, this factor has been largely overlooked in prior works, both in evaluations and in the development of mitigation strategies, underscoring our limited understanding of position bias. While Wang et al. (2023) explored the answer quality gap in score-based LLM-as-a-Judge evaluations using judgment scores, quantifying this variable in comparative settings (pairwise or list-wise) remains an open research problem. To address this gap, we propose the use of win rates (Zheng et al., 2024b) of candidate solutions against an expected baseline to measure the answer quality gap. This approach offers a practical quantitative metric for subsequent analysis and deepens our understanding of the factors driving position bias.
List-wise Evaluations: Although LLM-as-a-Judge has been categorized into pointwise, pairwise, and list-wise paradigms (Qin et al., 2024), most research has focused on pairwise comparative assessments due to their demonstrated effectiveness (Zheng et al., 2024b; Liusie et al., 2024). However, this leaves a significant gap in measuring and understanding the position bias in list-wise settings, where LLMs are prompted to evaluate three or more candidate solutions simultaneously. With advancements in model capabilities, list-wise evaluations present a potentially more cost-effective alternative to pairwise approaches, requiring fewer comparisons while maintaining consistency and fairness. Our work bridges this gap by extending position bias assessments from pairwise to list-wise settings, enabling a more comprehensive evaluation of LLM-as-a-Judge across varied configurations.

### v9

@@ Abstract
Lin Shi, Chiyu Ma, Wenhua Liang, Xingjian Diao, Weicheng Ma, Soroush Vosoughi Dartmouth College lin.shi.26@dartmouth.edu
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability in extensive evaluations zeng2023evaluating; karpinska2021perils. To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments zheng2024judging; li2024devbench; zhu2023judgelm, inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than evaluating content objectively, as shown in Fig. 1 (a). This bias has been observed across various types of LLM judges qin2024large; li2023split, raising concerns about their reliability. Prior studies have identified position bias alongside other biases and assessed its impact zheng2024large; zheng2024judging; zeng2023evaluating, but these investigations remain largely preliminary and lack a focused, systematic exploration. Although mitigation strategies have been proposed, they often suffer from incomplete bias removal guo2024bias, added complexity li2024split; khan2024debating; chua2024biasaugmented, the introduction of new biases ohi2024likelihood, inconsistent effectiveness gallegos2024bias, or impracticality for closed-source models requiring access to model internals wang2025eliminating; yu2025mitigate. Moreover, most empirical studies primarily examine position bias in pairwise settings wang2023large; zheng2024judging, leaving its underlying factors and behavior in more complex list-wise paradigms underexplored. These gaps underscore the need for a more systematic, in-depth exploratory analysis of position bias to better understand its origins, manifestations, and implications.
In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge. While evaluating with Position Consistency zheng2024judging, we further introduce two novel metrics: Preference Fairness and Repetition Stability. Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences. Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
@@ 2.2 Evaluation Metrics
Preference Fairness ($PF$) measures the extent to which LLM judges favor certain solution positions. In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference/bias for the first/second candidate model" used in previous studies zheng2024judging, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 1 (b). Previous studies proposed two common ways to measure the preference fairness. One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number (pcnpcn) and recency-count-number (rcnrcn). The counts are then normalized by the total number of prompt pairs (zheng2024judging; zhu2023judgelm). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like (li2023prd; liusie2024llm) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates (ipripr) and inconsistent recency rates (irrirr), where $ipr+irr=1$. However, this approach overlooks the fact that “position consistent judgments are also preference fair”, which leads to overly penalizing highly consistent LLM-judges.
@@
[…]
@@ 2.3 Factors Affecting Position Bias
Answer quality gap: While prior studies wang2023large explored quality disparities using "score gaps" in score-based LLM-as-a-Judge, this factor remains under-explored in comparative settings, which we address by introducing "answer quality gap" for both pairwise and list-wise evaluations. We define the quality of a candidate’s solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from one candidate model and the others to the same question and hence considered the Candidate-level factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs or series, it would prefer the highest-quality answer, where the corresponding candidate is denoted as the winner selected by the LLM judge.
Following this assumption, we measure the answer quality gap by the win rates of candidates over an expected baseline on a set of tasks and questions. However, if position bias occurs, the winner may be inconsistent when the order of candidate solutions is permuted in the query. Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen across all permutations (termed “consistent wins”), cases where there is no certain winner (termed “consistent ties”), and cases where different winners are selected after the solutions are permuted in the queries (termed “inconsistent judgment series"). We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment series ($C_{I}$), respectively. Inspired by zheng2024judging, we count inconsistent judgment pairs as ties for all candidate models, which is later calculated as a down-scaled win rate depending on the number of candidate models.
To calculate the win rates of candidate models for all three cases, we define the overall win rate (owrowr) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{p}(C_{t}+C_{I})}]$, where we have $p$ candidates in the list and $n$ judgment series. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-1/p|$, where $1/p$ is the expected baseline when all judgments are “ties”. In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment series) to quantify $\delta_{q}$ (zheng2024judging; li2023generative; raina2024llmasajudge), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook.
@@ 3.1 Experiment Settings
In this study, we evaluated position bias of 15 models from the GPT (openai2024gpt4), Claude (anthropic2024claude3), Gemini (geminiteam2024gemini), and Llama (Llama2023) series using our framework. For datasets, we adopted the modified MTBench (zheng2024judging) and DevBench (li2024devbench) due to their demonstrated high human-LLM agreement and the validated reliability of state-of-the-art LLM judges on the evaluation tasks. For pairwise comparisons, We fixed one of the candidates as vicuna-13b-v1.3 for MTBench and human for DevBench to serve as baselines, ensuring decent quality of solutions to the given questions. MTBench consists of 30 candidate models, 8 tasks, and 10 questions per task; for DevBench, we divide the general metric into more detailed ones and consider them as different tasks, resulting in 10 candidate models, 14 tasks, and 8 questions per task. We then paired solutions of these candidate models with that of the baseline candidate for evaluation by the LLM judges.
@@ 3.2 Empirical Results
The evaluation results of 12 close-source and 3 open-source LLM judges in terms of repetition stability, position consistency, and preference fairness on MTBench and DevBench are listed in Table 2. For each judge, we calculate its average $RS$, PCPC, and $PF$ across all candidates and tasks. For $RS$ and $PC$, higher values are preferable. A high $RS$ value is particularly important as a prerequisite for meaningful computations of $PC$ and $PF$, ensuring the LLM judge’s choice patterns are not due to random variations. Fig. 2 (a)(b) demonstrate that position bias varies by judges and tasks significantly. Fig. 2 (c) explores the correlation between the metrics $PC$ and $PF$. Fig. (d) to (f) further investigate the impact of the answer quality gap on position bias. These analyses were conducted by considering all judges together on MTBench. More analyses can be found in Appendix. Sec. D.
@@
RSRS PCPC PFPF Error RSRS PCPC PFPF Error PCPC PFPF Error
@@
Table 2: Evaluation results for Repetition Stability (RSRS), Position Consistency (PCPC), and Preference Fairness ($PF$) are presented for both pairwise and list-wise evaluation approaches, with the top 5 performances marked in bold. Errors arise from judgment failures (e.g., exceeding context window, not following output format). High error rates and low RSRS are marked red, rendering further evaluations invalid due to insufficient data. List-wise evaluation is conducted on four representative judges to validate scalability.
@@ 4 Main Findings
Figure 2: Judge performances on MTBench. Fig. (a)(b) are the radar charts for the PCPC comparison by family, judge, and task. Fig. (c) leverages linear regression to explore the general relationship between PCPC and $PF$. Fig. (d) to (f) investigate the impact of answer quality gap on position bias using overall win rates. Fig. (a) to (e) are for pairwise comparative settings, while Fig. (f) are obtained under list-wise evaluations.
@@
Position Bias Correlates to Answer Quality Gap: Intuitively, the difficulty of judging a pair of candidate answers is largely reflected by their difference in quality. In this study, as defined in Section 2.3, we quantify the quality gap ($\delta_{q}$) between candidate solutions and expected baseline (calculated by $1/p$ for a $p$-candidate list) by the overall win rate (owrowr). Therefore, $\delta_{q}$ increases as owrowr extends from baseline to 0 or 1. Fig. 2 (e) and (f) exhibit significant parabolic shapes, indicating that $PC$ is positively proportional to $\delta_{q}$. This aligns with our intuition that the answer pairs or series with larger quality disparities are easier to achieve judgment consistency, whereas those of similar quality are difficult to judge, increasing the likelihood of position bias that leads to lower $PC$. The same relationship is observed for each individual judge and across benchmarks, as demonstrated in Appendix.Sec. D.
@@ B.1 LLM-as-a-Judge
Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments (zheng2024judging). The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (chiang-lee-2023-large), adversarial attacks (chiang2023can), summarization (karpinska2021perils), machine translation (kocmi-federmann-2023-large), and instruction following (zeng2023evaluating), where models are tasked with scoring or ranking outputs. LLM-as-a-Judge paradigms range from pointwise, pairwise, and list-wise ranking approaches (qin2024large; zhu2024starlingb) to score-based and relation-based (comparative or ranking) settings (li2024split). Furthermore, explorations into multi-agent frameworks (li2023prd; chan2023chatevalbetterllmbasedevaluators) and meta-thinking or meta-judge systems (wu2024metarewardinglanguagemodelsselfimproving; li2025leveragingllmsmetajudgesmultiagent) have emerged to enhance reliability of LLM evaluations. Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness (zheng2024judging; ye2024justiceprejudicequantifyingbiases; ma2025judging), even in the most effective pairwise comparative settings.
@@ B.2 Position Bias
Position bias, the tendency of LLMs to favor certain positions within the prompt irrespective of actual content, is a pervasive issue observed across various domains and tasks (zheng2024judging; qin2024large; li2024devbench; li2024split). In the context of LLM-as-a-Judge, it typically refers to the position inconsistent scenarios where a LLM judge prefers different candidate solutions to a given question when the order of solutions in the prompt is permuted. To deal with such bias, the naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited. Then, researchers proposed ‘inconsistency-as-a-tie’ for both candidate models in pairwise comparative settings to consider all judgments for further analysis (zheng2024judging; li2023prd). This approach, while practically useful for evaluations, does not mitigate position bias.
Given the significance of position bias, researchers have developed debiasing strategies, including bootstrapping (hou2024large), split-and-merge techniques (li2023split), and multi-agent discussions (li2023prd; khan2024debating). While these methods demonstrate potential, they are often costly, time-consuming, or insufficient to eliminate position bias. More recently, mechanistic approaches have been explored to address position bias by modifying model internals like positional embeddings and causal attention mechanisms (yu2025mitigate; wang2025eliminating). These methods, while effective in eliminating position bias, lack applicability to state-of-the-art closed-source models, which are frequently used in practice but model internals are inaccessible. Therefore, position bias remains a complex and challenging issue to fully resolve.
B.3 Evaluating Position Bias Evaluation
Parallel to developing debiasing strategies, a crucial line of research focuses on comprehensively assessing position bias and identifying its influencing factors, thereby fostering a deeper understanding of this phenomenon. The evaluation of position bias typically concentrates on two main aspects: the consistency of judgments when the order of candidate solutions in the prompt is changed, and the preference direction (e.g., primacy or recency bias) when inconsistencies arise. Metrics for these aspects in pairwise comparisons are relatively established; however, their extension to, and rigorous measurement in, list-wise settings remain less explored.
While prior works (qin2024large; zhu2024starlingb) introduced list-wise ranking settings, they often become computationally expensive due to exponentially increasing number of pairs in the list or encounter significant degradation of performance due to ranking complexity. Moreover, the traditional preference direction measurement of position bias becomes increasingly complicated in list-wise paradigms (i.e., preferring the $i^{th}$ candidate solution for all i’s). However, practically, “choosing the best out of an n-item list” is a simpler but common need compared to exhaustive ranking of the entire list, which is not yet fully explored by prior works. A unified single metric for effectively measuring preference direction is therefore necessary to facilitate position bias evaluation in a more comprehensive manner.
B.4 Factors influencing Position Bias
Position bias may be influenced by several factors, including judge attributes, task types, candidate solution lengths and qualities, and more. For example, wang2023large discovered that judgment conflict rate (i.e., frequency where position bias exhibits across the dataset) is negatively correlated with the score gap between the two candidate responses. However, quantifying and understanding the impact of this “answer quality gap" in comparative settings (both pairwise and list-wise) is an area that requires further investigation.
Besides, prior works often observe position bias and conduct subsequent analyses based on single-instances judgments. However, observed biases might stem from random variations rather than systematic tendencies, hence compromising the reliability of downstream bias analyses. To mitigate this, researchers have employed multiple repetitions and used the mode judgment, an approach that, while robust, can be computationally expensive (chen2024mllmasajudge). In this sense, validating that position bias emerges from positional information rather than randomness becomes a crucial yet insufficiently-explored area.